OpenStudy (anonymous):

without graphing is the system independent dependent or inconsistent -2x-y=9, 3x-4y=-8 a. dependent b. inconsistent c. independent 2. without graphing is the system independent dependent or inconsistent 12x+3y=12, y=-4x+5

OpenStudy (cwrw238):

try solving the system

OpenStudy (cwrw238):

hint: first multiply the first equation by -4 then add

OpenStudy (cwrw238):

this gives: 8x + 4y = 36 3x - 4y = -8 now add

OpenStudy (cwrw238):

what does that give?

OpenStudy (cwrw238):

oops - i made a mistake! the first equation should be 8x + 4y = -36 3x - 4y = -8

OpenStudy (cwrw238):

r u there?